### v1

# ClustRecNet: A Novel End-to-End Deep Learning Framework for Clustering Algorithm Recommendation

Abstract
We introduce ClustRecNet – a novel deep learning (DL)-based recommendation framework for determining the most suitable clustering algorithms for a given dataset, addressing the long-standing challenge of clustering algorithm selection in unsupervised learning. To enable supervised learning in this context, we construct a comprehensive data repository comprising 34,000 synthetic datasets with diverse structural properties. Each of them was processed using 10 popular clustering algorithms. The resulting clusterings were assessed via the Adjusted Rand Index (ARI) to establish ground truth labels, used for training and evaluation of our DL model.
The proposed network architecture integrates convolutional, residual, and attention mechanisms to capture both local and global structural patterns from the input data. This design supports end-to-end training to learn compact representations of datasets and enables direct recommendation of the most suitable clustering algorithm, reducing reliance on handcrafted meta-features and traditional Cluster Validity Indices (CVIs). Comprehensive experiments across synthetic and real-world benchmarks demonstrate that our DL model consistently outperforms conventional CVIs (e.g. Silhouette, Calinski-Harabasz, Davies-Bouldin, and Dunn) as well as state-of-the-art AutoML clustering recommendation approaches (e.g. ML2DAC, AutoCluster, and AutoML4Clust). Notably, the proposed model achieves a 0.497 ARI improvement over the Calinski-Harabasz index on synthetic data and a 15.3% ARI gain over the best-performing AutoML approach on real-world data.
Codehttps://github.com/confanonymgit/clustrecnet
1 Introduction
Selecting an appropriate clustering algorithm for an unlabeled dataset remains one of the most elusive tasks in unsupervised learning. Due to the diversity of input data, numerous clustering algorithms have been proposed in the literature, each possessing its strengths and limitations. Several studies have been dedicated to the comparison and categorization of clustering algorithms (Brock et al. 2008; Arbelaitz et al. 2013; Rodriguez et al. 2019; Wani 2024), providing valuable insights into their advantages and limitations.
Despite these significant contributions, selecting the most suitable clustering algorithm for a given dataset remains a complex task, necessitating considerable domain knowledge and extensive trial-and-error experimentation. A crucial step in this process involves evaluating the quality of clustering solutions, which is typically carried out using established cluster validity indices (CVIs) (Hassan et al. 2024). Indices such as Silhouette (Sil) (Rousseeuw 1987), Calinski-Harabasz (CH) (Caliński and Harabasz 1974), Davies–Bouldin (DB) (Davies and Bouldin 1979), and Dunn (Dunn 1973) reflect several criteria of cluster compactness and separation. However, these CVIs often fall short when dealing with datasets characterized by complex high-dimensional structures, outliers, or overlapping clusters, thereby revealing some important limitations in existing cluster validation methodologies.
Recently, several AutoML-based approaches have been introduced to address the challenge of selecting the most suitable clustering algorithm for a given dataset in unsupervised settings (Francia, Giovanelli, and Golfarelli 2024). Many of them are inspired by established AutoML techniques in supervised learning, particularly those tackling the Combined Algorithm Selection and Hyperparameter optimization (CASH) problem (Thornton et al. 2013). The literature predominantly outlines two main strategies. The first of them jointly optimizes the algorithm choice and its hyperparameter space, as exemplified by AutoML4Clust (Tschechlov et al. 2021). However, such an approach is often computationally expensive, particularly on large and complex search spaces. The second, more prevalent, strategy treats the CASH problem sequentially — first selecting a suitable algorithm, followed by hyperparameter tuning. The approaches supporting this strategy leverage meta-learning techniques that transfer knowledge from existing datasets.
Early work by De Souto et al. (2008) pioneered the use of meta-learning for selecting the most appropriate clustering algorithm(s), thus laying the groundwork for subsequent research in this domain. Most AutoML systems for clustering follow a two-stage architecture: (1) Offline phase that extracts fixed-length meta-features and associates them with algorithm performance; (2) Online phase that employs shallow learners, such as $k$-nearest neighbors or decision trees, to recommend the most convenient clustering algorithms for new datasets. For example, AutoCluster (Liu, Li, and Tian 2021) leverages landmark-based and statistical meta-features, applying a grid search strategy for configuration optimization. In contrast, ML2DAC (Treder-Tschechlov et al. 2023) relies on a diversified set of data descriptors, including general properties (e.g. instance counts), statistical moments, and information-theoretic measures, to recommend both clustering algorithms and their optimal hyperparameters.
Despite these contributions, existing meta-learning-based methods face several inherent limitations. The reliance on fixed-length meta-feature vectors introduces an abstraction layer that may obscure essential structural and distributional characteristics of the data, which critically influence clustering behavior and algorithm performance. Moreover, the selection of informative and representative meta-features remains an open problem, with no established consensus on optimal descriptors (Poulakis, Doulkeridis, and Kyriazis 2024). Finally, the prevailing modular architecture – separating feature extraction from model training – limits the expressiveness and adaptability of learned representations, potentially constraining generalization to complex datasets.
In this paper, we propose a novel deep learning (DL) framework that eliminates the need for handcrafted meta-feature engineering or proxy representations. By treating each dataset holistically as a learning instance, we train a deep neural network in an end-to-end fashion to learn a direct mapping from raw data to algorithm recommendation. This design enables the model to capture high-level structural patterns directly from the underlying data distribution. As a result, it delivers more accurate and generalizable clustering algorithm recommendations across a diverse collection of synthetic and real-world datasets.
The paper is organized as follows: Section 2 briefly presents the data generation process, clustering algorithms being used, the proposed ClustRecNet model architecture, and the training process. Section 3 describes the experimental setup and discusses the results obtained from experiments with synthetic and real data. Finally, Section 4 concludes the study, summarizing its key findings and implications.
2 Methodology
The proposed recommendation system introduces a novel framework for selecting clustering algorithms, which overcomes key limitations of traditional validation indices and recent AutoML approaches. Our method employs a DL model trained on a diverse repository of 34,000 datasets, enabling it to learn structural patterns across varying clustering scenarios. To generate training labels, we evaluate the performance of 10 widely-used clustering algorithms on each dataset considered, using the Adjusted Rand Index (ARI) (Hubert and Arabie 1985) (i.e. the best-performing algorithm(s) correspond to the highest ARI value(s)). Once trained on these ARI-labeled datasets, our model can effectively recommend the most suitable clustering algorithms for previously unseen data. This design not only automates the selection process but also improves clustering accuracy and generalization, eliminating the need for handcrafted meta-features and CVIs.
This section provides a detailed exposition of the methodology used to develop the ClustRecNet framework. Here, we describe synthetic data employed for training and validation purposes, outline the clustering algorithms evaluated on these data, and present the details of the proposed network architecture along with the training procedure used to optimize its performance.
2.1 Data Generation
To train a robust DL-based recommendation framework, we generate synthetic datasets with controlled variation across key parameters, including the number and shape of clusters, the number of objects and features, intra-cluster density, degree of cluster overlap, and level of noise added to feature values. Accounting for these factors ensures that the model can generalize across a wide range of clustering scenarios and provide accurate algorithm recommendations.
Precisely, synthetic data were generated under two distinct scenarios:
1) Scenario 1 follows the data generation strategy proposed by Rodriguez et al. (2019). It relies on five key parameters: $K$ - the number of clusters in the dataset, $N$ - the number of objects, $D$ - the number of features (i.e. number of dimensions) for each object, $N_{e}$ - the number of objects per cluster, and $\alpha$ - the cluster separation parameter (i.e. degree of overlap between clusters). To cover a broad range of clustering complexities in our simulations, we varied the number of clusters, features, and objects per cluster. Specifically, we considered the values of 2, 5, and 10 for the number of clusters; 2, 5, 10, and 50 for the number of features (to simulate both low- and high-dimensional data); and 50 and 200 for the number of objects per cluster. This corresponds to 24 different parameter configurations. To select appropriate values of the $\alpha$ parameter, which controls inter-cluster separation and depends on both cluster count and dimensionality, we performed a grid search over the range [0.1, 10] with a step size of 0.1. The selected $\alpha$ values satisfied two criteria: No single clustering algorithm consistently achieved an ARI above 0.8 across all datasets of any given parameter configuration, and at least one algorithm exceeded this threshold in more than 20% of the realizations. These constraints prevented any single algorithm from clearly dominating the others. In total, we identified 11 parameter configurations meeting these criteria.
2) Scenario 2 corresponds to an alternative data generation strategy proposed by Zellinger and Bühlmann (2023). We used this strategy when no values of $\alpha$ were satisfying the desired ARI conditions for a given parameter configuration. Instead of considering the cluster separation parameter $\alpha$, the procedure of Zellinger and Buhlmann relies on a richer set of parameters, including cluster overlap, aspect ratio, radius ratio, cluster distribution, and imbalance ratio. In these settings, we explored three different minimum-maximum overlap intervals: $(0.001,0.002)$, $(0.001,0.25)$, and $(0.25,0.30)$. As the aspect ratio (i.e. the ratio between the longest and the shortest axis in a cluster), we used the values of 1 and 3. As the radius ratio (i.e. the ratio between the radii of the largest and the smallest clusters), we used the values of 1, 3, and 10. Cluster distributions considered included normal, exponential, and $t$ distributions. Finally, the cluster imbalance ratios of 1, 3, and 5 were explored. To mitigate bias, we calibrated parameter combinations under the same conditions as in Scenario 1. As a result, we identified 6 different valid parameter configurations.
Figure 1: Scatter plots illustrating the distribution of data points in two sample datasets: (a, Scenario 1) Clusters generated according to the strategy of Rodriguez et al. (2019), using the following parameters: $K=5$, $N=1000$, $D=2$, $N_{e}=200$, and $\alpha=5$; (b, Scenario 2) Clusters generated according to the strategy of Zellinger and Bühlmann (2023), using the following parameters: $K=10$, $N=500$, $D=2$, cluster overlap $(0.001,0.002)$, aspect ratio 33, radius ratio 33, normal distribution, and imbalance ratio 55.
Overall, we generated 2,000 different datasets for each of the 17 retained configurations (11 from Scenario 1 and 6 from Scenario 2) to ensure an empirical trade-off between dataset diversity and training stability, resulting in a comprehensive data repository of 34,000 synthetic datasets (available at our GitHub). Figure 1 illustrates examples of datasets considered in Scenarios 1 and 2. To ensure a consistent input scale for DL training, we normalized all datasets using $z$-scores. Additionally, we applied symmetric padding to align input shapes across varying dataset sizes.
2.2 Clustering Algorithms
The set of clustering algorithms considered in our study include partitioning methods such as $k$-means (McQueen 1967) and $k$-medians (Jain and Dubes 1988), hierarchical approaches such as Agglomerative Clustering (AC) (Defays 1977), BIRCH (Zhang, Ramakrishnan, and Livny 1996), and Ward’s method (Ward Jr 1963), and density-based techniques such as DBSCAN (Ester et al. 1996), HDBSCAN (Campello, Moulavi, and Sander 2013), and OPTICS (Ankerst et al. 1999). Moreover, we used the probabilistic Gaussian Mixture Models (GMM) (Reynolds et al. 2009) and the graph-based Spectral Clustering (SC) (Ng, Jordan, and Weiss 2001) clustering techniques for their unique properties and advantages. All implementations were carried out using the scikit-learn library, and where required (e.g. in $k$-means), the number of clusters was assumed to be known. Given the extensive literature on parameter selection for clustering algorithms, we did not analyze parameter effects comprehensively. Therefore, instead of optimizing hyperparameters for each realization, we performed grid search once per configuration to maximize the mean of ARI across all 2,000 datasets generated under each of our 17 parameter configurations, thus ensuring consistent and reproducible tuning, while significantly reducing computational overhead. The hyperparameter selection details for the 10 clustering algorithms considered in our study are available in Supplementary Material.
2.3 Deep Model Architecture
In our experiments, we comprehensively evaluated various neural network architectures to establish a robust baseline and ensure the efficacy of our proposed model. First, we examined feedforward neural networks (FNNs) with various complexities, implementing configurations with 3, 5, and 10 layers, to evaluate their performance on synthetic data from our repository. Subsequently, we explored basic CNN architectures incorporating 3 and 5 convolutional layers. These preliminary experiments provided critical insights into the strengths and limitations of various network architectures, supplying necessary information for our hybrid approach.
To extract the necessary features for clustering algorithm recommendation, we propose a novel DL architecture, integrating a CNN, two ResNets, and an attention mechanism. Such a hybrid model is engineered to harness the inherent advantages of each component, enhancing feature extraction, addressing the vanishing gradient issue, and highlighting crucial features within the input data. The selected architecture comprises four principal blocks discussed below. Its schematic representation is shown in Figure 2.
It starts with a convolutional layer, followed by batch normalization and max pooling. Integrating the 2D convolutional layer, batch normalization, and max pooling is a standard and effective way of organizing an initial processing block in contemporary CNN architectures. After the convolutional layer, the architecture incorporates two consecutive residual blocks. Each of them can be characterized by the following equation:
$\mathbf{Y}=F(\mathbf{X},\mathbf{W},\mathbf{b})+\mathbf{X},$ (1)
where $\mathbf{X}$ is the input to the layer, $F(\mathbf{X},\mathbf{W})$ is the residual mapping to be learned, $\mathbf{W}$ is the set of weight tensors in the residual block, and $\mathbf{b}$ is the bias term. Each residual block comprises two convolutional layers with batch normalization, employing the ReLU activation function. The residual connection, implemented through a shortcut, performs identity mapping or convolution for dimension matching, thereby bypassing the main convolutional layers of the residual block.
After the residual blocks, we implement a self-attention mechanism inspired by the transformer architecture from Vaswani et al. (2017). The attention block can be described by the following equation:
$Attention(\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{X})=softmax\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}+\mathbf{X},$ (2)
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are learned linear projections of the input features. Specifically $\mathbf{Q}=\text{Conv2D}(\mathbf{W}_{q},\mathbf{X}),\mathbf{K}=\text{Conv2D}(\mathbf{W}_{k}\mathbf{X})$, and $\mathbf{V}=\text{Conv2D}(\mathbf{W}_{v},\mathbf{X})$, where $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ are the respective learned weight matrices. The attention block uses the learned query, key, and value projections to compute a weighted sum of the input features, enabling the model to capture long-range dependencies within the feature maps and complement the local processing capabilities of the convolutional layers. The output from the attention block is subsequently flattened and processed through two fully connected layers, reaching the final layer that produces logits corresponding to the number of classes in the classification task.
Therefore, the proposed network architecture is designed to integrate local spatial pattern recognition (via CNNs), stable and hierarchical feature propagation (via residual connections), and global context modeling (via attention), offering a unified solution for learning discriminative representations from raw tabular data. As we will see in the next sections, an effective combination of these components allows our ClustRecNet model to deliver quality performance on complex clustering classification tasks.
Figure 2: Schematic representation of the proposed ClustRecNet model architecture.
2.4 Implementation Details
The described network architecture was chosen after an extensive experimentation with multiple network variants as it provided the best overall performance across our synthetic data repository. Moreover, the ablation study conducted on real-world data (see Section 3.2) showed that all components of the proposed network architecture are essential for generalizing across diverse clustering scenarios. Unlike traditional image-based CNNs, our model operates directly on tabular data encoded as 2D arrays, where rows correspond to individual data points (objects) and columns to features. We treat each dataset as a single-channel ”image” of shape $(1,N,D)$, where $N$ is the number of objects (padded to a fixed height), and $D$ is the number of features. This tabular treatment allows the convolutional network to learn local feature patterns across both object and feature dimensions.
The network starts with a 2D convolutional layer that takes a single-channel input and produces a single-channel output. It uses a 3×3 kernel with a stride of 1 and padding of 1 to preserve spatial dimensions, followed by a batch normalization and a 2×2 max pooling operation to reduce resolution and improve generalization. Next, two residual blocks extract hierarchical features. The first residual block expands the number of channels from 1 to 2 and reduces spatial dimensions by using a stride of 2 in the first convolutional layer. The second block further increases the channels from 2 to 3, using a stride of 2 as well. Each residual block contains two 3×3 convolutional layers, followed by batch normalization. If the spatial or channel dimensions change, a shortcut path with a 1×1 convolution and batch normalization ensures proper residual addition.
Figure 3: Normalized confusion matrices for 10 clustering algorithms generated by the proposed model on synthetic test data.
To capture long-range interactions, we incorporated a self-attention block into the model architecture, which follows the residual layers. Query and key projections reduce the channel dimension using 1×1 convolutions (with a reduction factor of 8), while the value projection maintains the input channels. Attention weights are computed across flattened spatial positions and applied back to the input with a residual connection. The output is flattened and passed through two fully connected layers - the first compresses the representation to 16 neurons, while the second produces predictions over clustering algorithms. The final model was trained using the Adam optimizer (learning rate $10^{-4}$, weight decay $7\times 10^{-3}$) with the Binary Cross-Entropy (BCE) loss with logits. Training proceeds by 30 epochs with a batch size of 32, using 10-fold cross-validation. All experiments were conducted on an NVIDIA A100-SXM4-40GB GPU. All implementation details, including source code and configuration, are available at our GitHub.
Table 1: Comparison of F1-score, Hamming distance, and ARI results across different methods (ClustRecNet and Silhouette, CH, DB, and Dunn CVIs) on the synthetic test data. The results are reported as means $\pm$ standard deviations.
Method F1-score ($\uparrow$) Hamming ($\downarrow$) ARI ($\uparrow$)
ClustRecNet 0.757 $\pm$ 0.017 0.183 $\pm$ 0.011 0.878 $\pm$ 0.007
Silhouette 0.720 $\pm$ 0.009 0.280 $\pm$ 0.009 0.365 $\pm$ 0.020
CH 0.669 $\pm$ 0.013 0.331 $\pm$ 0.013 0.381 $\pm$ 0.017
DB 0.681 $\pm$ 0.014 0.319 $\pm$ 0.014 0.344 $\pm$ 0.016
Dunn 0.637 $\pm$ 0.011 0.363 $\pm$ 0.011 0.280 $\pm$ 0.017
3 Results and Discussion
3.1 Experiments with Synthetic Data
In this section, we rigorously evaluate ClustRecNet that uses as input an unlabeled $objects\times features$ dataset and outputs the most suitable algorithm(s) for its clustering, among those described in Section 2.2. During training, we used 90% of the data for 10-fold cross-validation to ensure robustness and reduce overfitting, with the remaining 10% reserved for final model validation. Additionally, weight decay regularization was applied through the Adam optimizer to further prevent model overparameterization and enhance generalization performance.
The BCE loss consistently decreased, indicating effective learning, from an initial point of 0.528 to 0.360 for training, and from 0.457 to 0.370 for validation. This reduction translates into F1-score improvements: For training, the F1-score increased from 0.534 to 0.763, while for validation it increased from 0.635 to 0.752, reflecting the model’s growing prediction accuracy. The Hamming distance decreased from 0.295 to 0.177 for training, and from 0.249 to 0.186 for validation, marking, respectively, a 40% and 25.3% reduction in misclassification.
To evaluate the effectiveness of our recommendation framework, we compared its performance on the test data against the following popular CVIs: Silhouette, CH, DB, and Dunn. We performed a grid search over predefined thresholds for each CVI to determine the one that yields the highest F1-score when compared to binary labels derived from ARI. The best selected threshold was then used to choose the most suitable clustering algorithm. The obtained F1-score, Hamming distance, and ARI results are reported in Table 1.
To compute the ARI, the clustering algorithm, selected either by our model or by a CVI, was executed using the hyperparameter configurations obtained via grid search, as detailed in Supplementary Material. Our model provided a mean F1-score of 0.757 and a mean ARI of 0.878, outperforming the leading CVIs (Silhouette and CH) by more than 5% in terms of F1-score and by 0.497 in terms of ARI. The mean Hamming distance was reduced by 53% compared to Silhouette, confirming the superiority of our method in predicting suitable clustering algorithms. We also carried out the Wilcoxon signed-rank test to assess whether the improvements achieved by our proposed model over traditional CVIs were statistically significant. The test results confirm the superiority of our model, with the obtained $p$-values less than 0.005 across all three evaluation metrics, strongly rejecting the null hypothesis of equal performance.
To provide a more detailed breakdown of prediction outcomes across 10 clustering algorithms considered in our study, we present their confusion matrices (see Figure 3). All reported values are normalized, making them directly comparable across algorithms. As shown, our proposed model consistently identifies the most suitable algorithms with high precision and recall, particularly using Spectral Clustering, Agglomerative Clustering, DBSCAN, HDBSCAN, and OPTICS, each exhibiting an accuracy over 80%.
Furthermore, we conducted experiments with various random sampling techniques to mitigate the effect of data imbalance. Despite our efforts, these approaches failed to provide significant result improvements, indicating the need for more sophisticated novel approaches that can effectively handle the nuances of multi-label imbalanced datasets.
3.2 Experiments with Real Data
Components Median Mean
All: CNN + ResNet + ATN (CH) 0.1775 0.2295
All: CNN + ResNet + ATN (Sil) 0.1836 0.2279
No CNN: ResNet + ATN (CH) 0.1229 0.2232
No CNN: ResNet + ATN (Sil) 0.1129 0.1808
No ResNet: CNN + ATN (CH) 0.0955 0.1719
No ResNet: CNN + ATN (Sil) 0.1129 0.1756
No ATN: CNN + ResNet (CH) 0.0797 0.2062
No ATN: CNN + ResNet (Sil) 0.0890 0.1522
Table 2: Results of our ablation study on real-world data presented in terms of ARI.
To assess the effectiveness of the proposed recommendation framework in practical settings, we conducted experiments on 10 well-known real-world datasets from the UCI Machine Learning Repository [^1] https://archive.ics.uci.edu. These datasets, summarized in Table 3, exhibit a diverse range of characteristics, including the number of objects, features, and classes, encompassing a variety of domains and complexities. Since all these datasets include ground-truth labels, we were able to compute ARI for each clustering solution obtained to compare the performance of our recommendation model to those of state-of-the-art AutoML-based clustering methods, including ML2DAC, AutoCluster, and AutoML4Clust (AML4C).
For each real-world dataset considered, our model was first applied to predict the most suitable clustering algorithm. The selected algorithm was then executed with default scikit-learn parameters over a predefined range of candidate cluster counts. Either Silhouette or CH index was then used to determine the optimal number of clusters. Such an approach simulates a typical unsupervised clustering scenario, where neither ground-truth labels nor the number of clusters are available at inference time.
After selecting a clustering algorithm and estimating the number of clusters, we performed a grid search to identify suitable hyperparameters for that specific algorithm. The grid values for each parameter were predefined based on commonly used configurations in the literature to ensure consistency across our evaluation. The selected clustering algorithm was then used with the best configuration found, while the clustering quality was assessed using ARI.
To isolate the contribution of each architectural component of our model, we performed an ablation study in which one of its main components - either the CNN block, or two ResNet blocks, or the attention mechanism (ATN) - was removed, one at a time. The results, summarized in Table 2, highlight the importance of all three main components working in synergy. The full CH-based model provided the highest mean ARI of 0.2295 and a median ARI of 0.1775. In contrast, the average ARI across all ablated models dropped to 0.1850 (for mean) and to 0.1021 (for median), corresponding to losses of 21.5% and 73.8%, respectively. Compared to the closest CH-based ablation, the improvement in mean was rather modest, i.e. 2.8%, but the median gain was 44.4%, suggesting a much more stable clustering performance. Similarly, the full Silhouette-based model outperformed the best ablation by 26.1% in terms of the mean ARI and by 62.6% in terms of the median ARI, confirming the cumulative value of all model’s components.
To conduct the clustering performance comparison on real-world data, we first considered a baseline CNN using the first CNN block of our model architecture. This baseline shared all other elements of the pipeline – including training data, cluster estimation, and evaluation setup, allowing for a fair performance comparison. We further benchmarked our model against state-of-the-art AutoML-based clustering recommendation systems. Specifically, we carried out the official, containerized implementation of ML2DAC (available on GitHub [^2] https://github.com/tschechlovdev/ml2dac) – endorsed with a reproducibility badge by SIGMOD – on the same real-world data. Furthermore, we evaluated two built-in meta-learning approaches from the same container: AutoCluster and AML4C, covering a range of meta-learning strategies. A comprehensive summary of all comparative results is presented in Table 3.
Dataset Objects Features Classes ClustRecNet Baseline CNN AutoCluster AML4C ML2DAC
CH Sil CH Sil DB
Breast tissue 106 9 6 0.2682 0.2682 0.2682 0.2398 0.2398 0.2398 0.2099 0.2099
Ecoli 336 7 8 0.5011 0.5011 0.5011 0.5011 0.5011 0.5011 0.1535 0.3863
Glass 214 9 7 0.2453 0.2453 0.2453 0.1399 0.1399 0.1399 0.0994 0.1472
Haberman 306 3 2 0.1040 0.1040 0.0274 0.1556 0.0621 0.0621 0.0323 0.0323
Iris 150 4 3 0.5681 0.5681 0.5681 0.5584 0.5438 0.5438 0.5681 0.5681
Parkinsons 195 22 2 0.0514 0.1218 0.0514 0.0000 0.0000 0.0000 0.0473 0.0473
Transfusion 748 4 2 0.0193 0.0193 0.0131 0.0174 0.0304 0.0095 0.0196 0.0344
Vehicle 846 18 4 0.1097 0.0869 0.0878 0.0125 0.0119 0.0088 0.1380 0.1333
Vertebral column 310 6 3 0.3669 0.3669 0.1361 0.3669 0.0967 0.3369 0.1211 0.2885
Wine quality red 1599 11 6 0.0611 -0.0029 0.0157 0.0000 0.0000 0.0000 0.0505 0.0505
Median 0.1775 0.1836 0.1119 0.1477 0.0794 0.1010 0.1102 0.1403
Mean 0.2295 0.2279 0.1914 0.1991 0.1626 0.1842 0.1440 0.1898
Table 3: Detailed comparison of clustering algorithm recommendation methods across 10 real-world datasets in terms of ARI.
Comparison $p$-value
ClustRecNet (CH) vs Baseline CNN 0.0312
ClustRecNet (CH) vs AutoCluster (CH) 0.0391
ClustRecNet (CH) vs AutoCluster (Sil) 0.0039
ClustRecNet (CH) vs AutoCluster (DB) 0.0020
ClustRecNet (CH) vs AML4C 0.0195
ClustRecNet (CH) vs ML2DAC 0.0371
Table 4: Wilcoxon signed-rank test $p$-values reflecting the performance improvement provided the proposed ClustRecNet model (CH variant) over the basic CNN and AutoML approaches on real-world data.
Moreover, we analyzed the ARI score distribution using a boxplot diagram (see Figure 4) and carried out the Wilcoxon signed-rank test to assess the performance improvement provided by ClustRecNet over the other competing methods (see Table 4). We found that the new model demonstrated statistically significant ARI improvements, with the Wilcoxon test $p$-values << 0.05, over the state-of-the-art AutoML-based clustering frameworks, including AutoCluster (CH, Silhouette, and DB variants), AML4C, and ML2DAC, as well as over the baseline CNN. Overall, our model systematically yielded the best results in terms of the mean and median ARI, regardless the cluster validity index (CH or Silhouette) used to determine the number of clusters, highlighting the model’s stability across different CVIs.
While the proposed model demonstrated strong performance and stability, several limitations remain. First, although the model generalizes well to real-world datasets, its predictive accuracy depends on the diversity and coverage of synthetic data used for training. Expanding the training set to better reflect edge-case scenarios, such as highly imbalanced clusters or very noisy data, could further improve its robustness. Second, our current pipeline selects the number of clusters using internal validation indices (CH or Silhouette in our experiments), while all of them are known to be imperfect. Incorporating into it a learned cluster count estimator or an ensemble of estimators may offer more reliable performance in the future. Third, the model assumes a tabular input representation; extending it to accommodate graph-based or time-series clustering problems could broaden its applicability.
Figure 4: Boxplot comparison of clustering algorithm recommendation methods across 10 real-world datasets.
4 Conclusion
We developed ClustRecNet – a novel recommendation framework to identify the most suitable clustering algorithms for a given dataset, showing that deep learning models can be effectively used to tackle this challenging issue. First, we generated a variety of synthetic datasets, employing different data generation strategies. Second, we applied 10 popular clustering algorithms on each of the datasets to evaluate their performance. The suitability of clustering algorithms for each dataset was assessed using ARI and encoded as a binary vector, convenient for multi-label classification tasks. Third, we designed and tested a DL-based hybrid network architecture with a CNN block, two residual blocks, and an attention mechanism. Such an architecture allows us to extract both low-level and high-level data features, with the attention mechanism highlighting critical features.
The proposed recommendation framework based on this architecture was evaluated across a variety of synthetic and real-world benchmarks, consistently outperforming conventional cluster validity indices, such as Silhouette, Calinski-Harabasz, Davies-Bouldin, and Dunn, as well as state-of-the-art AutoML approaches, such as ML2DAC, AutoCluster, and AutoML4Clust, thus offering an enhanced practical solution for unsupervised learning tasks.
Future research could build upon this framework in several directions. First, integrating into the model pipeline dimensionality reduction techniques, such as principal component analysis or autoencoders, could further optimize the system’s performance in high-dimensional spaces, particularly in scenarios where the feature set is large or sparsely populated. Additionally, exploring alternative feature extraction methods, such as graph neural networks, may enhance the model’s ability to capture complex relationships within data, especially for datasets with inherent hierarchical or graph-based structure. Another promising avenue is extending our framework to optimize the parameter settings for each algorithm recommended, possibly using reinforcement learning, thus widening its applicability and enhancing precision.